OpenStudy (pphalke):

Suppose the APR on a credit card is 19%. If a card carries an average daily balance of $600.00 over one billing period of 30 days, what is the finance charge for that period?

OpenStudy (pphalke):

@jim_thompson5910

OpenStudy (austinl):

Considering 14.9 is the average APR one thing you have is bad credit...

OpenStudy (austinl):

Also how much are you paying in that 30 day period?

OpenStudy (pphalke):

I have no idea

OpenStudy (austinl):

If you pay the entire ammount, than your finance charge will be around $9.58.

jimthompson5910 (jim_thompson5910):

finance charge = (ADB)*(APR)*(time) ADB is the average daily balance APR is the annual percentage rate time is in years. If you're given 1 month, then time is 1/12 years

OpenStudy (pphalke):

it needs a monthly payment amount and this question doesn't have that

OpenStudy (pphalke):

@austinl

OpenStudy (austinl):

In order to find finance charges you need a time peiod.

OpenStudy (pphalke):

30 days

OpenStudy (pphalke):

@austinl

OpenStudy (pphalke):

@jim_thompson5910 please help, I got the idea how do I solve?

jimthompson5910 (jim_thompson5910):

finance charge = (ADB)*(APR)*(time) ADB = 600 APR = 0.19 time = 30/365

OpenStudy (pphalke):

I get 3420 is it right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's way too big

OpenStudy (pphalke):

whats the answer? I don't know what to calculate

jimthompson5910 (jim_thompson5910):

what is 30/365 equal to?